OpenStudy (anonymous):

how do u find the maximum for r=4sin2(theta)? is sq2/2 wrong?

OpenStudy (anonymous):

r=4 is the max..

OpenStudy (anonymous):

does the 2 in front of the theta not affect the maximum in any way?

OpenStudy (anonymous):

any sinusoidal function has max of 1 so r max=4*1=4

OpenStudy (anonymous):

the two will not affect in the max. value of r but it will affect the value for which the max. occurs.. if it is 4sin (theta) then theta=pi/2 for maxima..but if it is r=4sin (a theta) then theta= pi/(2a)

OpenStudy (anonymous):

thanks! what if there's a number in front of the sin theta like in r=3-5sin2theta? how would that change the maximum?

OpenStudy (anonymous):

r will be maximum when 5 sin 2theta is minimum...because it is subtracted from 3... so minimum of sin x is -1 so rmax=3-5*(-1)=3+5=8 sin 2theta=-1 theta=-pi/4

OpenStudy (anonymous):

sorry i meant r=3-5cos2theta but the minimum of cosx is also -1 so the final answer would still be the same right?